# Diagnosing Quantum Many-body Chaos in Non-Hermitian Quantum Spin Chain via Krylov Complexity

### v1

@@ Abstract
Krylov variance and reciprocity.— Beyond the growth of $C_{K}$, a metric known as Krylov variance [22] has been proposed to measure the fluctuation of the Lanczos coefficients, which characterizes the localization of the Krylov wave function as an analogue to Anderson localization [48]. In Hermitian systems, the Krylov variance is defined as $\sigma_{K}^{2}={\rm Var}\{\ln|b_{2n-1}/b_{2n}|\}$ [22]. Similarly, we define the variance of the non-Hermitian system by replacing $b_{n}$ with $j_{n}=|b_{n}c_{n}|^{1/2}$, yielding
$\sigma_{K}^{2}={\rm Var}\left\{\ln\frac{j_{2n-1}}{j_{2n}}\right\}\quad,\quad n=1,2,3,\cdots.$ (4)
@@
Figure 2(c) shows the Krylov variance first grows with increasing $W_{\gamma}$ and saturates at larger $W_{\gamma}$. Through finite-size scaling, we find $\sigma_{K}^{2}$ can be universally fitted with function $y=\sigma_{K}^{2}L^{-\beta}$ against $x=(W_{\gamma}-W_{L})L^{\alpha}$ [see inset of Fig. 2(c)]. This allows us to identify a critical point, $W_{L}=0.0351$, with exponents $\alpha=1.816$ and $\beta=0.112$. When $W_{\gamma}<W_{L}$, the Krylov variance is small, which results in the delocalization of the Krylov wave function [Fig. 1(d)]. When $W_{\gamma}>W_{L}$, the Krylov wave function localizes in a finite region. Here, the system is still chaotic, as supported by the level repulsion (as discussed below). Yet, the localized Krylov wave function and hence the suppression of Krylov complexity emerge as a signature of ergodicity breaking [21, 29, 33].
@@
In non-Hermitian systems, Lanczos coefficients $b_{n}$ and $c_{n}$ are not complex conjugates, violating the reciprocity of the Krylov space. To quantify this, we evaluate arguments of the Lanczos coefficients, $\theta_{n}=\arg(b_{n}c_{n})$, such that $\cos\theta_{n}=1$ represents reciprocal tunneling, and $\cos\theta_{n}=-1$ gives the maximally non-reciprocal tunneling, which is associated with rapid relaxation. Figures 2(b1-b4) show the profiles of $\cos\theta_{n}$ as $W_{\gamma}$ increases. Specifically, for $W_{\gamma}<W_{L}$, the entire Krylov space remains reciprocal, supporting delocalization of the Krylov wave function. For $W_{L}<W_{\gamma}<W_{C}$, tunneling within $n<L$ is reciprocal, but a rapid relaxation occurs as the Krylov wave function spreads to $n>L$. For $W_{\gamma}>W_{C}$, the Krylov wave function rapidly relaxes within $n<L$, as the first few sites become maximally non-reciprocal.
@@
To leverage the sign-flip behavior of $\cos\theta_{n}$, we introduce a parameterized metric,
@@
which is insensitive to the choice of $d$ for a certain range smaller than $L$. We show this reciprocal metric in Fig. 2(d) for $d=4$, and more examples in SM. For different $L$, $R_{K}^{(d)}$ collapses to $y=R_{K}^{(d)}$ with $x=(W_{\gamma}-W_{C})L^{\alpha}$, yielding the mean critical point $W_{C}=1.647$ and exponent $\alpha=0.820$ for $d\in\{4,5,6\}$, with the relative deviation $<1\%$. The universal scaling thus shows the transition from the chaotic to non-chaotic phases.
@@
Entanglement entropy and level spacing.— To provide a comprehensive view of the aforementioned phase transitions, we bridge the findings from the Krylov complexity with conventional signatures of quantum chaos including entanglement entropy and complex level spacing. It is known that the chaos-non-chaos transition can be signified by the divergence of the variance of the bipartite entanglement entropy [40]. We show the standard deviation of the bipartite entanglement entropy, $\sigma_{S}$, in Fig. 3(a). A divergent peak is observed as $L\to\infty$. The data can be universally described by function $y=\sigma_{S}L^{-\beta}$ with respect to $x=(W_{\gamma}-W^{\prime}_{C})L^{\alpha}$ [inset of Fig. 3(a)]. This leads to the critical point $W^{\prime}_{C}=1.763$ with exponents $\alpha=2.013$ and $\beta=1.228$. Furthermore, a critical behavior at very small $W_{\gamma}$ can be seen in Fig. 3(b). Using a similar scaling method, we identify the critical point $W^{\prime}_{L}=0.0409$ and exponents $\alpha=1.514$ and $\beta=-0.054$. The critical points extracted from $\sigma_{S}$ are similar to the ones of the Krylov complexity. In addition, we obtain a different critical point at $W^{*}=3.236$ with the finite-size analysis of the entanglement entropy, $S$ (see SM). This attributes to the transition to integrability as the system is well described by product states when $W_{\gamma}$ is sufficiently large.

### v2

@@ IV Krylov variance and reciprocity
In order to quantitatively determine the values of $W_{L}$ and $W_{C}$, we turn to the Lanczos coefficients. Beyond the growth of $C_{K}$, a metric known as Krylov variance [24] has been proposed to measure the fluctuation of the Lanczos coefficients, which characterizes the localization of the wave function as an analogue to Anderson localization [74]. In Hermitian systems, the Krylov variance is defined as $\sigma_{K}^{2}={\rm Var}\{\ln|b_{2n-1}/b_{2n}|\}$ [24]. Similarly, we define the variance of the non-Hermitian system by replacing $b_{n}$ with $j_{n}=|b_{n}c_{n}|^{1/2}$, yielding
$\sigma_{K}^{2}={\rm Var}\left\{\ln\frac{j_{2n-1}}{j_{2n}}\right\},\quad n=1,2,3,\cdots.$ (5)
@@
Figure 4(a) shows that $\sigma_{K}^{2}$ first grows with increasing $W_{\gamma}$ and saturates at larger $W_{\gamma}$. Through finite-size scaling, we find that $\sigma_{K}^{2}$ can be universally fitted with function $y=\sigma_{K}^{2}L^{-\beta}$ against $x=(W_{\gamma}-W_{L})L^{\alpha}$ [see Fig. 4(b)]. This allows us to identify a critical point, $W_{L}=0.0351$, with exponents $\alpha=1.816$ and $\beta=0.112$. When $W_{\gamma}<W_{L}$, the variance is small, which results in the delocalization of the wave function, $\varphi_{n}$ [Fig. 1(d)]. When $W_{\gamma}>W_{L}$, $\varphi_{n}$ localizes in a finite region. Here, the system is still chaotic, as supported by the level repulsion (as discussed below). Yet, the localized wave function and hence the suppression of complexity emerge as a signature of ergodicity breaking [23, 34, 38].
@@
In non-Hermitian systems, Lanczos coefficients $b_{n}$ and $c_{n}$ are not complex conjugates, violating the reciprocity of the Krylov space. To quantify this, we evaluate arguments of the Lanczos coefficients, $\theta_{n}=\arg(b_{n}c_{n})$, such that $\cos\theta_{n}=1$ represents reciprocal tunneling, and $\cos\theta_{n}=-1$ gives the maximally nonreciprocal tunneling, which is associated with rapid relaxation. Figure 3 shows the profiles of $\cos\theta_{n}$ as $W_{\gamma}$ increases. Specifically, for $W_{\gamma}<W_{L}$ [Fig. 3(a)], the entire Krylov space remains reciprocal as $\braket{\cos\theta_{n}}\approx 1$ for all $n$, supporting complete delocalization of $\varphi_{n}$. For $W_{L}<W_{\gamma}<W_{C}$ [Figs. 3(b) and 3(c)], tunneling within $n<L$ is reciprocal as $\braket{\cos\theta_{n}}\approx 1$, but a rapid relaxation occurs as $\varphi_{n}$ spreads to $n>L$ with $\braket{\cos\theta_{n}}\approx 0.5$. For $W_{\gamma}>W_{C}$ [Fig. 3(d)], $\varphi_{n}$ rapidly relaxes within $n<L$, as the first few sites become maximally nonreciprocal with $\braket{\cos\theta_{n}}\approx 0$. This significantly inhibits the spreading of Krylov wave function and the complexity growth, much more influential than the fluctuations of the Lanczos coefficients.
Figure 4: (a), (b) The localization to nonlocalization transition predicted by the Krylov variance $\sigma_{K}^{2}$, which exhibits a steep increase at the critical point, $W_{L}=0.0351$, determined by the finite-size scaling analysis (b). (c), (d) Chaotic-nonchaotic transition determined by the Krylov reciprocity $R_{k}^{(4)}$, which shows a sign flip at the critical point $W_{C}=1.647$. The result is consistent with the finite-size scaling (d).
@@
To leverage the sign-flip behavior of $\cos\theta_{n}$, we introduce a parametrized metric,
@@
which is insensitive to the choice of $d$ for a certain range smaller than $L$. We show this reciprocal metric in Fig. 4(c) for $d=4$, and more examples in Appendix E. For different $L$, $R_{K}^{(d)}$ collapses to $y=R_{K}^{(d)}$ with $x=(W_{\gamma}-W_{C})L^{\alpha}$ [Fig. 4(d)], yielding the mean critical point $W_{C}=1.647$ and exponent $\alpha=0.820$ for $d\in\{4,5,6\}$, with the relative deviation $1\%$. The proposed metrics from scaling analysis can signify the phase transitions linked to quantum chaos, and the phase diagram involving the two critical points, $W_{L}$ and $W_{C}$, is qualitatively illustrated in Fig. 1(c).
V Complex level spacing and entanglement entropy
Figure 5: (a), (b) Mean radial distribution of complex level spacing ratio (CSR), $\langle r\rangle$, for different ranges of $W_{\gamma}$. Dot lines (red) are the extrapolated values when $L\to\infty$. See examples shown in the inset of panel (a), and details in Ref. [75]. Blue arrows indicate the phase transition points obtained from Krylov reciprocity for $W_{C}$ and Krylov variance for $W_{L}$, respectively. (c), (d) Mean angular distribution of CSR, $\langle\cos\theta\rangle$, for different ranges of $W_{\gamma}$. (e) Distributions of complex level spacing ratios for $W_{\gamma}=0.005,\ 0.8,$ and 5.0\ 5.0 and $L=12$ (from left to right), corresponding to GOE, AI class, and 2D Poisson ensemble, respectively.
To provide a comprehensive view of the aforementioned phase transitions, we bridge the findings from the Krylov complexity with conventional signatures of quantum chaos including complex level spacing and entanglement entropy.
It is widely recognized to determine quantum chaos by the complex level spacing ratio (CSR), $r_{j}=(z_{j,{\rm NN}}-z_{j})/(z_{j,{\rm NNN}}-z_{j})$, where $z_{j}$ is an eigenvalue of Hamiltonian (1), and $z_{j,{\rm NN}}$ and $z_{j,{\rm NNN}}$ are its nearest-neighbor and next-nearest-neighbor eigenvalues [12]. The signatures of CSR are represented by its mean radial distribution $\langle r\rangle=\int|r|P(r){\rm d}r$ as shown in Figs. 5(a) and 5(b), and the angular distribution by $\langle\cos\theta\rangle=\int\cos\theta P(r){\rm d}r$ ($\theta=\arg r$) shown in Figs. 5(c) and 5(d). Increasing $W_{\gamma}$, the level statistics change from the Gaussian orthogonal ensemble (GOE) to $\rm AI$$^{\dagger}$ class, and then to the 2D Poisson ensemble [Fig. 5(e)]. To minimize the finite-size effects, we extrapolate $\langle r\rangle_{\infty}$ in the limit $L\to\infty$ [red dots in Figs. 5(a) and 5(b)] [75]. The values of $\langle\cos\theta\rangle$ do not show very clear scaling relations as $\langle r\rangle$ does, but also exhibit similar behavior to $\langle r\rangle$ as it transits between different values for different phases, and qualitatively reflects the spreading of the eigenvalues from the real axis to the whole complex plane.
As $W_{\gamma}\to 0$, $\langle r\rangle_{\infty}=0.567$, consistent with $\langle r\rangle_{\rm GOE}=0.5689$ [11]. The angular signature is also close to $\langle\cos\theta\rangle_{\rm GOE}=-0.4038(2)$ [76]. A plateau appears when $0.025\lesssim W_{\gamma}\lesssim 1.6$, where $\langle r\rangle_{\infty}=0.726(3)$, close to $\left\langle r\right\rangle_{\rm AI^{\dagger}}=0.7222$ [14], and the angular signature $\langle\cos\theta\rangle$ also touches $\langle\cos\theta\rangle_{\rm AI^{\dagger}}=-0.1727(2)$ for large $L$. This indicates that the system belongs to the $\rm AI$$^{\dagger}$ class. Increasing $W_{\gamma}$ greater than $W_{C}$, the level statistics deviate from the AI class and become the 2D Poisson ensemble. In this region, $\langle r\rangle_{\infty}=0.6673(4)$, close to the 2D Poisson ensemble where $\left\langle r\right\rangle_{\rm Poi(2D)}=2/3$ and $\left\langle\cos\theta\right\rangle_{\rm Poi(2D)}=0$ [12]. Therefore, the statistical signatures of CSR are consistent with the results of Krylov complexity.
Figure 6: (a) and (b) Standard deviation of bipartite entanglement entropy, $\sigma_{S}$, for different ranges of $W_{\gamma}$. The finite-size scaling [insets in (a) and (b)] yields critical points $W_{C}^{\prime}=1.763$ and $W_{L}^{\prime}=0.0409$.
Another notable signature of phase transition is the bipartite entanglement entropy [53]. We show the standard deviation of the bipartite entanglement entropy $\sigma_{S}$, in Fig. 6(a). A divergent peak is observed as $L\to\infty$. The data can be universally described by function $y=\sigma_{S}L^{-\beta}$ with respect to $x=(W_{\gamma}-W^{\prime}_{C})L^{\alpha}$ [inset of Fig. 6(a)]. This leads to the critical point $W^{\prime}_{C}=1.763$ with exponents $\alpha=2.013$ and $\beta=1.228$. Furthermore, a critical behavior at very small $W_{\gamma}$ can be seen in Fig. 6(b). Using a similar scaling method, we identify the critical point $W^{\prime}_{L}=0.0409$ and exponents $\alpha=1.514$ and $\beta=-0.054$. The critical points extracted from $\sigma_{S}$ are similar to the ones of the Krylov complexity. In addition, we obtain a different critical point at $W^{*}=3.236$ with the finite-size analysis of the entanglement entropy $S$ (see Appendix F). This attributes to the transition to integrability as the system is well described by product states when $W_{\gamma}$ is sufficiently large.
VI Conclusions
In conclusion, we have studied non-Hermitian quantum chaos in the disordered XY model, revealing two critical phase transitions, including the chaotic-nonchaotic transition and Krylov localization. Through examining the Krylov complexity and Lanczos coefficients characterized by their variance and reciprocity, we identify the critical points and exponents of these phase transitions. Our results are corroborated by established measures including entanglement entropy and complex level spacings, which showcase the applicability of the complexity and associated measures in the study of quantum chaos. The sign flip of the reciprocity provides a quantitatively more precise criterion for chaos suppression than the entanglement entropy. Furthermore, we observe intricate chaotic behavior in non-Hermitian systems, where prethermalization may exist in the Krylov delocalized regime.
The non-Hermitian spin model can be realized in Rydberg atom arrays, where spin states are encoded by two Rydberg states [77]. The XY interactions correspond to the dipole-dipole interactions between the Rydberg states, and the dissipation can be induced by coupling the Rydberg states to low-lying electronic states [78]. Randomness of the local fields can be introduced by ac Stark shifts with addressing lasers. In generic quantum simulators, non-Hermiticity can be realized through Sz.-Nagy dilation [79, 80] and linear combination of unitaries [81, 82, 66]. Finally, to measure the Krylov complexity, it has been proposed to apply measurable bases for estimation [83], and the measure of Loschmidt echo is potentially related to the ergodicity breaking [Fig. 1(d)].
Acknowledgements.
We are grateful to Dong-Ling Deng, Dong Yuan, Si Jiang, Chang Liu (SQZI), Chang Liu (NUS), Tianyi Yan, Xiaopeng Li, Xingze Qiu, and Saud Čindrak for fruitful discussions. Y.Z. acknowledges support from the Shanghai Qi Zhi Institute Innovation Program SQZ202317. W.X. acknowledges support from the National Natural Science Foundation of China/Hong Kong RGC Collaborative Research Scheme (Project CRS CUHK401/22) and New Cornerstone Science Foundation. L.L. acknowledges support from the National Key Research and Development Program of China (Grant No. 2021YFA1402003) and the National Natural Science Foundation of China (Grants No. 12374329 and No. U21A6006). W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1, the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26), and the International Research Collaboration Fund of the University of Nottingham.
Appendix A Krylov space and bi-Lanczos algorithm
Input: $H$, $\psi_{0}$
Output: Lanczos coefficients: $\{a_{n}\}$, $\{b_{n}\}$, $\{c_{n}\}$; Krylov basis: $\{p_{n}\}$, $\{q_{n}\}$
$b_{0},c_{0}\leftarrow 0$,  $p_{0},q_{0}\leftarrow\psi_{0}$
// Initialize
1 for $n=0$ to d2d-2 do // $d=2^{L}$ is the dimension of the Hilbert space and $L$ is the system size
2 $p_{n+1},\ q_{n+1}\leftarrow Hp_{n},\ H^{\dagger}q_{n}$
3 $a_{n}\leftarrow q_{n}^{\dagger}p_{n+1}$
 $p_{n+1}\leftarrow p_{n+1}-a_{n}p_{n}-c_{n}p_{n-1}$
 // $p_{-1}=0$
 $q_{n+1}\leftarrow q_{n+1}-a_{n}^{*}q_{n}-b_{n}^{*}q_{n-1}$
 // $q_{-1}=0$
 $W\leftarrow[p_{0},p_{1},\cdots,p_{n}][q_{0},q_{1},\cdots,q_{n}]^{\dagger}$
 // Complete reorthogonalization
4 $\text{res}\leftarrow 0$
5 while $\text{res}<0.707$ do
6    $\tilde{p}_{n+1}\leftarrow p_{n+1}-Wp_{n+1}$
7    $\tilde{q}_{n+1}\leftarrow q_{n+1}-W^{\dagger}q_{n+1}$
8    $\text{res}\leftarrow\min\left(\lVert\tilde{p}_{n+1}\rVert/\lVert p_{n+1}\rVert,\lVert\tilde{q}_{n+1}\rVert/\lVert q_{n+1}\rVert\right)$
9    $p_{n+1},\ q_{n+1}\leftarrow\tilde{p}_{n+1},\ \tilde{q}_{n+1}$
10
11   end while
 $b_{n+1}\leftarrow\lVert p_{n+1}\rVert$
 // The argument of $b_{n}$ is fixed to be 0.
12 $c_{n+1}\leftarrow q_{n+1}^{\dagger}p_{n+1}/b_{n+1}$
13 $p_{n+1},\ q_{n+1}\leftarrow b_{n+1}^{-1}p_{n+1},\ c_{n+1}^{*-1}q_{n+1}$
14
15 end for
@@ 13 $p_{n+1},\ q_{n+1}\leftarrow b_{n+1}^{-1}p_{n+1},\ c_{n+1}^{*-1}q_{n+1}$
Algorithm 1 Bi-Lanczos Algorithm with Complete Reorthogonalization
The study of Krylov complexity has its roots in solving Schrödinger equations, which starts with the temporal evolution of the wave function,
$\psi\left(t\right)=e^{-iHt}\psi_{0}=\sum_{n}{\frac{\left(-it\right)^{n}}{n!}H^{n}\psi_{0}}.$ (7)
In this context, the Hamiltonian $H$ and the initial wave function $\psi_{0}$ spans a Krylov space,
$\mathcal{K}(H,\psi_{0})=\mathrm{span}\{\psi_{0},H\psi_{0},H^{2}\psi_{0},\cdots\}.$ (8)
Contrasting with the Hermitian scenario, a non-Hermitian Hamiltonian generates a bi-orthogonal system, which includes an additional dual space,
$\mathcal{K}^{*}(H,\psi_{0})=\mathrm{span}\{\psi_{0},H^{\dagger}\psi_{0},H^{\dagger}{}^{2}\psi_{0},\cdots\}.$ (9)
These spaces can be orthogonalized to yield two sets of vectors, $\mathrm{span}\{p_{0},p_{1},p_{2},\cdots\}$ and $\mathrm{span}\{q_{0},q_{1},q_{2},\cdots\}$, which fulfill the bi-orthogonality condition,
$q_{m}^{\dagger}p_{n}=\delta_{mn}.$ (10)
It is noteworthy that the corresponding matrices $P=[p_{0},p_{1},p_{2},\cdots]$ and $Q=[q_{0},q_{1},q_{2},\cdots]$ are not unitary. Consequently, the Hamiltonian can be tridiagonalized through the relation,
$T=Q^{\dagger}HP,$ (11)
with $Q^{\dagger}P=PQ^{\dagger}=I$ given that $P$ and $Q$ are both square and invertible.
This process can be numerically executed via the bi-Lanczos (two-sided Lanczos) iteration. Here, we summarize the basic idea of bi-Lanczos algorithm. First, we assume that we have finished $n$ iterations, and obtained biorthonormal basis $\mathcal{K}_{n}=\{p_{0},p_{1},\cdots,p_{n}\}$ and $\mathcal{K}_{n}^{*}=\{q_{0},q_{1},\cdots,q_{n}\}$. Apart from Eq. (10), we can further let
$q_{m}^{\dagger}q_{n}=\delta_{mn},\quad p_{m}^{\dagger}p_{n}\neq\delta_{mn},$ (12)
because only 2 out of 3 orthogonality conditions can be simultaneously satisfied in bilinear systems.
Then, we can find the next basis $p_{n+1}$ by hitting $p_{n}$ with $H$, and removing its projection remained in $\mathcal{K}_{n}$. This procedure is similar to Gram-Schmidt elimination process. The projector is constructed by the bases in the dual space $\mathcal{K}^{*}_{n}$, and finally $Hp_{n}$ can be written as
$Hp_{n}$ $=b_{n+1}p_{n+1}+\left(\sum_{m=0}^{n}q_{m}q_{m}^{\dagger}\right)Hp_{n},$
$=b_{n+1}p_{n+1}+\sum_{m=0}^{n}\left(q_{m}^{\dagger}Hp_{n}\right)q_{m},$ (13)
where $b_{n+1}$ is a normalization factor to be determined. Now, we consider the summation part. If m<n1m<n-1, $q_{m}^{\dagger}Hp_{n}=\left(H^{\dagger}q_{m}\right)^{\dagger}p_{n}=0$, because $H^{\dagger}q_{m}\in\mathcal{K}^{*}_{m+1}$, $p_{n}\in\mathcal{K}_{n}$ but $p_{n}\notin\mathcal{K}_{m+1}$. Therefore, the summation over $m$ can be reduced to m=n1m=n-1 and $n$ only, and we eventually have a three-term recurrence relation
$Hp_{n}=b_{n+1}p_{n+1}+a_{n}p_{n}+c_{n}p_{n-1},$ (14)
with $a_{n}=q_{n}^{\dagger}Hp_{n}$ and $c_{n}=q_{n-1}^{\dagger}Hp_{n}$. The coefficients $b_{n}$, $a_{n}$, and $c_{n}$ are the diagonal and sub-diagonal elements of the tridiagonal matrix $T$ in Eq. (11). This recurrence relation can also be represented as $HP=PT$, and with the help of Eq. (11), we can also obtain the recurrence relation for $q_{n}$ which reads
$H^{\dagger}q_{n}=c^{*}_{n+1}q_{n+1}+a^{*}_{n}p_{n}+b^{*}_{n}q_{n-1}.$ (15)
To circumvent the breakdown issue, we have employed the complete reorthogonalization algorithm in this study. The full pseudo-code of this algorithm is shown in Algorithm 1. We note that there are also several more expeditious methods, such as selective and partial reorthogonalization [84, 85, 73].
With this methodology, the amplitude of the wave function within the Krylov space is defined by,
$\varphi_{n}(t)=q_{n}^{\dagger}\psi(t).$ (16)
The temporal evolution of the coefficients $\varphi_{n}(t)$, akin to an effective tight-binding model, reads
$i\frac{\partial}{\partial t}\varphi_{n}=b_{n}\varphi_{n-1}+a_{n}\varphi_{n}+c_{n+1}\varphi_{n+1}.$ (17)
We note that the Arnoldi iteration, also known as the modified Gram-Schmidt algorithm, can deal with more general forms of $H$ with good numerical stability, albeit resulting in a more intricate recurrence relation, where the tridiagonal matrix $T$ is replaced by a Hessenberg matrix.
Appendix B Krylov wave functions
Figure S1: Krylov wave function at intermediate time $t=300$ (blue) and long time $t=1000$ (red) for different $W_{\gamma}$ and fixed $L=10$.
Figure S1 presents several examples of the Krylov wave function for varying $W_{\gamma}$ at times $t=300$ and $1000$, providing additional data depicted in Fig. 1(c). Our analysis here corroborates that the wave function exhibits localization around $n\approx 0$ for $W_{\gamma}<W_{L}$ ($W_{L}\approx 0.0351$) on the intermediate time scale ($t=300$), while delocalization is observed in the long term ($t=1000$). The apparent flatness of the delocalized wave functions is attributed to the logarithmic scaling of the $x$-axis, with significant occupation occurring at large $n$ values.
When $W_{\gamma}$ is comparable or larger than $W_{L}$, the Krylov wave function rapidly reaches a steady state characterized by localized features. The difference between the wave functions at $t=300$ and $t=1000$ are invisible. It is noted that the central peak shifts to smaller $n$ values as $W_{\gamma}$ increases.
When $W_{\gamma}>W_{C}$ ($W_{C}=1.647$), the saturated Krylov complexity remains relatively constant. The critical behavior may not be directly reflected in the Krylov wave function, however, as elaborated in the main text, it can be inferred from the Lanczos coefficients. We introduce the concept of Krylov reciprocity to identify the critical point, which provides a deeper understanding of the system at the phase transition.
Figure S2: Krylov inverse participation at intermediate time $t=300$ (dashed) and long time $t=1000$ (solid) for different $W_{\gamma}$ and system size $L=6,8,10$.
To quantitatively evaluate the localization of the Krylov wave function, we show the Krylov inverse participation ratio (IPR) in Fig. S2 for $t=300$ and $t=1000$, respectively. With higher values of Krylov IPR, the Krylov wave function is more localized. At $t=300$, the Krylov IPR is very large for $W_{\gamma}<10^{-2}$, which corresponds to a prethermal behavior where the Krylov wave function is localized around $n\approx 0$. At $t=1000$, the Krylov IPR monotonically increases as $W_{\gamma}$ increases, indicating that the Krylov wave function becomes more localized. There are several plateaus for Krylov IPR, and the ramping regimes correspond to the critical points of the phase transitions at $W_{L}$ and $W_{C}$.
Appendix C Krylov complexity growth for different system size and at early times
Figure S3 shows extended results for Krylov complexity growth across system sizes L=4, 6, 8L=4,\ 6,\ 8, complementing the results in Fig. 2(a). The observed dynamics closely match those presented in the main text. When $W_{\gamma}<W_{L}$, we consistently observe a plateau in $C_{K}(t)$ as $L$ increases, indicating prethermal behavior. A linear envelope applied to the mid-term growth for $W_{L}<W_{\gamma}<W_{C}$ across all system sizes. Furthermore, the long-time saturation values of $C_{K}(t)$ follow systematic scaling relationships with $L$, which are shown in Fig. 2(b,c).
Figure S3: Growth of Krylov complexity for (a) $L=4$, (b) $L=6$ and (c) $L=8$. The dahsed lines are scaling of $\propto t^{2}$ and dashed-dot lines for $\propto t$.
The early-time growth of the Krylov complexity, $C_{K}$, exhibits quadratic behaviors. We fit $C_{K}$ using the function $C_{K}(t)=at^{2}$ for $t<0.01$, and present the coefficients for different system sizes $L=4,6,8,10$ and a range of $W_{\gamma}$ from 0.01 to 10 in Fig. S4. Our analysis reveals that the coefficient $a$ is proportional to $L$ and exhibits a parabolic relationship with $W_{\gamma}$, approximated as $a\simeq L(1+0.33W_{\gamma}^{2})$.
Figure S4: Early-time growth rate of the Krylov complexity, $a=\lim_{t\to 0}C_{K}/t^{2}$. Solid lines are the fitting with function $a=L(1+\frac{1}{3}W_{\gamma}^{2})$. The fitting errors are smaller than sizes of the scatters.
Appendix D Fitting of Lanczos coefficients with Tsallis $q$-log statistics
Figure S5: Fitting of the Lanczos coefficients with Tsallis $q$-log function.
Reference [46] proposes using the Tsallis $q$-log function to fit the Lanczos coefficients for large $n$, with $|b_{n}|^{2}\propto-\ln_{q}(n/2^{L})$. The $q$-log function is defined as